OpenStudy (anonymous):

Graph the function f(x) = –3x3 + 9x2 – 2x + 3 using graphing technology and describe the end behavior. I Just need the 2nd part of the question

OpenStudy (anonymous):

as \(x\to -\infty\) you have \(f(x)\to +\infty\) and as \(x\to \infty\) you have \(f(x)\to -\infty\) you can see it from the graph